# Low energy phonons in single crystal ZrW2O8

### v1

@@ III.2 DFT calculations
Fig. 11 shows the PDOS with neutron weighting applied (see eq. 1 in the supplementary information [32]), which allows direct comparison with the neutron structure factor measured in an experiment, such as that in ref. 24. The scattering lengths of all three constituent elements are similar in magnitude, as are their Debye-Waller factors, so the largest effect on the neutron cross-section comes from the division of all the other terms by the atomic mass. This gives significantly more weight to the oxygen modes than to tungsten and zirconium, with the consequence that the neutron scattering signal is dominated by oxygen motion. We can use a similar procedure for the single crystal plots shown in figs. 24, i.e. of element-wise decomposition of the signal. Examples are shown in fig. 12, for the $(-2,2,2)$ Brillouin zone center, i.e. equivalent to fig. 3. As with the PDOS, most of the signal arises from oxygen motions, with contributions from tungsten and zirconium around an order of magnitude smaller. However, although the neutron measurements are not directly sensitive to the motion of the heavier metal atoms, the fact that the DFT gives good agreement for the oxygen means that we can be confident that its predictions for the zirconium and tungsten motion are accurate, since each of them are surrounded by a cage of oxygen.
IV Discussion
From the results and analysis presented so far, we can see that there is generally very good agreement between the phonon dispersion relations and structure factors computed from DFT, and our inelastic neutron scattering measurements. Of course, anharmonicity is essential to describe any kind of thermal expansion, negative or positive. We have already cited extensively the predictions made by Gupta et al [30], which were based on a phenomenological theory developed for understanding librational modes at $q=0$ in the $\alpha$ phase of solid nitrogen [49]. In this approach the phonon modes are decomposed into the displacement and rotation of molecular units (analogous to the RUM approach) and the computed potential rewritten as a polynomial to $4^{\rm{th}}$ order in terms of these coordinates. This is clearly an approximation, though it should be acknowledged that it does a reasonable job, estimating correctly that in ZrW2O8 at some wavevectors the effects of implicit and explicit anharmonicity seem to approximately cancel each other.
It would seem, then, that for the low energy excitations studied here there is a cancellation of two large effects on the energy of the modes. Both arise due to anharmonicity, and given the unconventional behavior of ZrW2O8 it is clear that anharmonicity has an important role in governing the material’s thermal properties. In order to cancel the mode softening expected from the mode Grüneisen parameters, large amplitude thermal motions are required. This would naturally arise in a structure that is especially flexible, with individual atoms subject either only to weak constraints or strongly constrained only in one or two directions.
Alternative methods exist in addition to DFT of computing the phonon spectra, that include anharmonicity. These include the temperature-dependent effective potential (TDEP) method [50, 51, 52] and the self-consistent phonon (SCPH) method [53, 54]. These would in principle allow us to investigate in detail the effect of temperature and anharmonicity on the phonon modes, and hence NTE. However, the calculations to do this for a system like ZrW2O8 would be prohibitively computationally expensive, and are thus outside the scope of the present work.
Notwithstanding the aforementioned shortcomings of the DFT computations, it should be acknowledged that they perform an admirable job of reproducing many of the details of the INS measurements, both in terms of the eigenvalues (mode energies) and eigenvectors (measured neutron scattering intensities). This then gives high confidence that they can be used to give both a qualitative and quantitative commentary on the nature of the atomic displacements involved in the low energy phonon modes measured here. This was discussed extensively already by Gupta et al [30], who correlated the translational and rotational motions of quasi-rigid ZrO6 octahedra and WO4 tetrahedra, together with distortions of those polyhedra, with reported distortions found from local structure probes [55]. We can also have good confidence that the DFT results can be used as a general interpretative tool, e.g. in the flexibility analysis of NTE undertaken by Rimmer et al [28].
Molecular dynamics simulations [29] suggest that the oxygens that bridge Zr–O–W bonds are subject to large transverse vibrations. The distortions we observe from our DFT computations indicate that some modes do indeed include such motions. Because of the relative rigidity of the WO4 tetrahedra (it is well-known that the W–O bonds are shorter than the Zr–O bonds, so the WO4 tetrahedra are more tightly bonded and stiffer than the ZrO6 octahedra [56]) a consequence of such vibrations is that the lone unbonded oxygen on each tetrahedron is also subject to large vibrations [25]. Indeed, displacements of the lone oxygen would have a lower energy cost due to weaker bonding. The flexibility analysis performed by Rimmer et al [28], using DFT results similar to those reported here, showed that the low energy phonon modes responsible for most of the NTE (those with negative Grüneisen parameters) had largely fixed Zr–O and W–O bond lengths. These results together suggest that the tension effect description of NTE offers an appropriate description of the behavior of ZrW2O8 : rigid Zr–O and W–O bonds together with a large transverse vibration of oxygens that bridge Zr and W lead naturally to a reduced distance between Zr and W and hence NTE. By validating the DFT that underpins this interpretation against INS data, we are able to give greater credence to it. We also note that in the high pressure orthorhombic phase of $\gamma$-ZrW2O8 [57] the lone oxygen becomes bonded to a tungsten ion, making its motion more constrained than in the ambient pressure cubic $\alpha$-ZrW2O8 phase, which is then manifest in a significantly smaller (negative) coefficient of thermal expansion in the high pressure polymorph.
V Conclusion
In conclusion, we present a comprehensive set of INS measurements of the phonons in ZrW2O8 , and compare them to predictions from DFT. We find that we have to compress the lattice compared to the equilibrium in the DFT in order to get agreement with the neutron data, but that after doing so the agreement is excellent. This gives confidence that the overall picture of atomic motions predicted by DFT closely matches the experimental situation, so that DFT can be confidently used as a basis for additional analysis of NTE. The calculated mode Grüneisen parameters suggest that most low energy phonon modes should soften on warming as the lattice shrinks, but instead most modes either remain unchanged or harden slightly, as well as broadening somewhat. This can be understood in terms of anharmonicity arising from an increased amplitude of atomic motions associated with the phonon modes.
Acknowledgements.
We are grateful to D. Voneshen (ISIS) and R. Fair (STFC-SCD) for their advice and assistance on the use of Euphonic. We are indebted to M. Dove (QMUL) for his insightful and informative comments. Raw data are available via ref. [58]. We thank the Institut Laue-Langevin for the allocation of beamtime for the experiments reported here. Computing resources were provided by the STFC Scientific Computing Department’s SCARF cluster.

### v2

@@ Acknowledgements.
We can decompose the simulations shown in figs. 2 – 4 of the main text, into the contributions from the different atomic species. Examples are shown in fig. 13, for the $(-2,2,2)$ Brillouin zone center, i.e. equivalent to fig. 3 from the main text. Most of the signal arises from oxygen motions, with contributions from tungsten and zirconium around an order of magnitude smaller. A similar effect is seen for computation of the phonon density of states. However, although the neutron measurements are less directly sensitive to the motion of the heavier metal atoms, the fact that the DFT gives good agreement for the oxygen means that we can be confident that its predictions for the zirconium and tungsten motion are accurate, since each of them are surrounded by a cage of oxygen ions.
Appendix IV: Structure factors from MD calculations
Here we show the neutron structure factor $S(\mathbf{q},\omega)$ computed from the MD calculations at several fixed values of $\mathbf{q}$, for temperatures of 100, 200, 300 and 400 K. The peaks were fitted using Gaussians, and it is the center positions of these that are plotted in fig. 10 of the main text. As well as the peaks moving to higher energies with increasing temperature, one can also see that, as would be expected, they also broaden.
Figure 14: Energy scans of the MD-computed neutron structure factor $S(\mathbf{q},\omega)$ at fixed wavevectors $\mathbf{q}=(0,0.5,0)$ (blue circles) , $\mathbf{q}=(0.25,0.5,0)$ (green circles), and $\mathbf{q}=(0.5,0.5,0)$ (red circles). Data are successively offset on the y-scale by 2 units. Fits of the peaks to a Gaussian lineshape are shown by black lines. These data are from the MD simulation at $T=100$ K.
Figure 15: Energy scans of the MD-computed neutron structure factor $S(\mathbf{q},\omega)$ at fixed wavevectors $\mathbf{q}=(0,0.5,0)$ (blue circles) , $\mathbf{q}=(0.25,0.5,0)$ (green circles), and $\mathbf{q}=(0.5,0.5,0)$ (red circles). Data are successively offset on the y-scale by 2 units. Fits of the peaks to a Gaussian lineshape are shown by black lines. These data are from the MD simulation at $T=200$ K.
Figure 16: Energy scans of the MD-computed neutron structure factor $S(\mathbf{q},\omega)$ at fixed wavevectors $\mathbf{q}=(0,0.5,0)$ (blue circles) , $\mathbf{q}=(0.25,0.5,0)$ (green circles), and $\mathbf{q}=(0.5,0.5,0)$ (red circles). Data are successively offset on the y-scale by 2 units. Fits of the peaks to a Gaussian lineshape are shown by black lines. These data are from the MD simulation at $T=300$ K.
Figure 17: Energy scans of the MD-computed neutron structure factor $S(\mathbf{q},\omega)$ at fixed wavevectors $\mathbf{q}=(0,0.5,0)$ (blue circles) , $\mathbf{q}=(0.25,0.5,0)$ (green circles), and $\mathbf{q}=(0.5,0.5,0)$ (red circles). Data are successively offset on the y-scale by 2 units. Fits of the peaks to a Gaussian lineshape are shown by black lines. These data are from the MD simulation at $T=400$ K.